# PEL-NAS: Search Space Partitioned Architecture Prompt Co-Evolutionary LLM-driven Hardware-Aware Neural Architecture Search

### v2

@@ Appendix A Algorithm
Algorithm 1 provides a detailed, step-by-step description of the PEL-NAS framework. The process begins with a one-time training of a zero-cost (ZC) ensemble predictor. The core of the algorithm is a parallel evolutionary search conducted independently within several disjoint niches ($\mathcal{S}_{k}$), which are defined by architectural complexity. In each generation, an LLM acts as an intelligent evolutionary operator to generate a new candidate architecture ($A_{child}$) under the niche-specific constraints. The candidate is then evaluated using the pre-trained predictor and direct hardware lookup, and the Pareto archive for that niche ($\mathcal{P}_{k}$) is updated. Finally, all niche archives are aggregated and filtered through a non-dominated sort to produce the final, comprehensive Pareto front.
Algorithm 1 PEL-NAS: Partitioned Evolutionary LLM-driven NAS
@@ Appendix B Result of Ablation Study on all Datasets and Devices
This section provides a comprehensive visualization of the ablation studies discussed in the main paper’s Section 4. We present the full set of Pareto fronts for each of the three datasets—CIFAR-10, CIFAR-100, and ImageNet16-120—across all six hardware devices from the HW-NAS-Bench benchmark. These figures visually supplement the aggregated quantitative results presented in Table 5 and demonstrate the consistent and crucial contribution of each component within the PEL-NAS framework.
In each subplot, the reader can clearly observe that the Pareto front discovered by the full PEL-NAS model (in blue) consistently envelops and dominates the fronts from the three ablated versions. This provides strong visual evidence that each key component of our framework—the partitioning strategy, the LLM operator, and the ZC ensemble predictor—is critical for discovering the optimal trade-off between accuracy and latency across diverse datasets and hardware constraints.
Figure 6: Results of the ablation study on CIFAR-10 across six hardware devices. Each subplot compares the Pareto fronts discovered by our full model (PEL-NAS) against its three ablated versions. The consistent dominance of the full PEL-NAS model demonstrates that each component is crucial for discovering the optimal trade-off between accuracy and latency
@@ Appendix E LLM Prompt Templates and Co-evolution Process
This appendix provides the full prompt structures used in the two stages of each PEL-NAS generation and explains how they form a tight co-evolution loop.

### v3

@@ Appendix A Algorithm
Algorithm 1 provides a detailed, step-by-step description of the LLM-NAS framework. The process begins with a one-time training of a zero-cost (ZC) ensemble predictor. The core of the algorithm is a parallel evolutionary search conducted independently within several disjoint niches ($\mathcal{S}_{k}$), which are defined by architectural complexity. In each generation, an LLM acts as an intelligent evolutionary operator to generate a new candidate architecture ($A_{child}$) under the niche-specific constraints. The candidate is then evaluated using the pre-trained predictor and direct hardware lookup, and the Pareto archive for that niche ($\mathcal{P}_{k}$) is updated. Finally, all niche archives are aggregated and filtered through a non-dominated sort to produce the final, comprehensive Pareto front.
Algorithm 1 LLM-NAS: Partitioned Evolutionary LLM-driven NAS
@@ Appendix B Result of Ablation Study on all Datasets and Devices
This section provides a comprehensive visualization of the ablation studies discussed in the main paper’s Section 4. We present the full set of Pareto fronts for each of the three datasets—CIFAR-10, CIFAR-100, and ImageNet16-120—across all six hardware devices from the HW-NAS-Bench benchmark. These figures visually supplement the aggregated quantitative results presented in Table 5 and demonstrate the consistent and crucial contribution of each component within the LLM-NAS framework.
In each subplot, the reader can clearly observe that the Pareto front discovered by the full LLM-NAS model (in blue) consistently envelops and dominates the fronts from the three ablated versions. This provides strong visual evidence that each key component of our framework—the partitioning strategy, the LLM operator, and the ZC ensemble predictor—is critical for discovering the optimal trade-off between accuracy and latency across diverse datasets and hardware constraints.
Figure 6: Results of the ablation study on CIFAR-10 across six hardware devices. Each subplot compares the Pareto fronts discovered by our full model (LLM-NAS) against its three ablated versions. The consistent dominance of the full LLM-NAS model demonstrates that each component is crucial for discovering the optimal trade-off between accuracy and latency
@@ Appendix E LLM Prompt Templates and Co-evolution Process
This appendix provides the full prompt structures used in the two stages of each LLM-NAS generation and explains how they form a tight co-evolution loop.